OpenStudy (anonymous):

Simplify: X^2+13x+40 ------------ X^2+6x-16 A. X+5 ----- X+2 B. X+5 ----- X-2 C. X+5 ----- X+8 D. X+8 ----- X-2 E. X-5 ----- X-2

OpenStudy (anonymous):

you need to factor them out as in (X+a number)*(X+a number) To do this you need to find two numbers that when you add the give you 13 and when you multiply they give you 40

OpenStudy (anonymous):

then the same to the bottom to that add to 6 and multiply to -16 hint one of the numbers will be negative

OpenStudy (anonymous):

its 6x and 16 so it would be 6x minus 16

OpenStudy (anonymous):

and then 13x and 40 13x plus 40

OpenStudy (anonymous):

not quite, forget about the x's the formula is (x+a)(x+b)= x^2+(a+b)x+ab so what we are looking at is the (a+b)=13 and A*B= 40

OpenStudy (anonymous):

Okay I think I get what you're saying now

OpenStudy (anonymous):

The denominator is the same A+B= 6 A*B= -16 what numbers did you find?

OpenStudy (anonymous):

Well for the 13 and 40 one, I found that 8 plus 5 equaled 13 and 5x8 equaled 40

OpenStudy (anonymous):

@garrett_payne

OpenStudy (anonymous):

8 and -2 kinda tricky

OpenStudy (anonymous):

Yeah I couldn't figure out what it was

OpenStudy (anonymous):

So what would I do ext @garrett_payne

OpenStudy (anonymous):

next*

OpenStudy (anonymous):

what do you have now?

OpenStudy (anonymous):

(x+5)(x+8)= x^2+(8+-2)x+8(-2) ??

OpenStudy (anonymous):

\[((X+8)*(X+5))/ (X+8)*(X-2)\] do you see anything in common? Similar terms cancel out

OpenStudy (anonymous):

Cancel out the 8

OpenStudy (anonymous):

kinda, cancel out (x+8) you are left with (X+5)/(X-2)

OpenStudy (anonymous):

Yeah thats what I meant

OpenStudy (anonymous):

So it would be B right?

OpenStudy (anonymous):

(x+5)(x+8)= x^2+(8+-2)x+8(-2) This is not correct as the denominator should be in the (X+a)(X+b) form like you did in the numerator

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

there you go, done Yes, B

OpenStudy (anonymous):

Cool Thanks :)